OpenStudy (anonymous):

I have no clue here please help.. f(x)=8−6sinπ5(x+1)Find the first negative value of x for which f(x) = 3

OpenStudy (amistre64):

let f(x) = 3 and algebra thru it

OpenStudy (amistre64):

f(x)=8−6sinπ5(x+1) 3=8−6sinπ5(x+1) -5=−6sinπ5(x+1) 5/6= sinπ5(x+1) i got no idea what the rest of that is spose to mean

OpenStudy (amistre64):

\[ maybe:~\sin~[5π(x+1)]\] if so then inverse arcsin(5/6) = 5π(x+1) arcsin(5/6)/5π = x+1 arcsin(5/6)/5π - 1 = x

OpenStudy (anonymous):

do you have any idea what the answer would be

OpenStudy (anonymous):

I've been working on this for close to an hour and im no where near getting it

OpenStudy (amistre64):

apart from inputing that into a calculator ... no

OpenStudy (amistre64):

using a derivative, i know that at -.9 there is the first miminum point using the calculations above i get about -.93729, which is to far the difference between them is .03729 -.90000 .03729 -------- - .86271 gets rather close